Reports 1-1 of 1 Clear search Modify search
terrence.tsang - 15:46 Thursday 25 November 2021 (18996) Print this report
SR2 IM L stuck?

I was trying the measure transfer functions for all DoFs for SR2 but I got a strange response in the IM L direction.
Fig.1 shows the transfer function for IM L, and it seems very low and there's no peak, whereas it looks fine for other DoFs, see Fig.2 for IM T for example.
There's no coherence between IM L actuation and all readouts, including the TM optical levers.

I tried putting -20000/20000 counts (which is for some reason the limit) IM L actuation. Looking at ndscope IM L didn't move and stayed at around 568.69 um with really small RMS.
The raw H1 OSEM readout is very close to zero, which may be a sign of a disconnected/broken sensor?
Looking at TM L, it didn't seem to respond when I put actuation offset in IM L as well. But I am not sure if this is because it's too weak or if the coil isn't working.
There's a chance that both sensing and actuation are somehow malfunctioning, or it's touching somewhere.

Images attached to this report
Comments to this report:
terrence.tsang - 16:11 Thursday 25 November 2021 (18999) Print this report

Yokozawa-san posted something about IM H1 cable disconnected in 18687 and Fabian plugged it back in today. Fig.1 shows a healthy IM L transfer function.

Images attached to this comment
Search Help